10) Find 5-1 (22 + 2x + 2)3 2+1 dar.
Answer #1

sol:- Given LL at) (1?+23+2) 3 da Let - u= x 72x72 ; du - 204+2 +0 वर i du=2(x+1) de ; (+1) dx = du u=610²+21-1 +2 = 1 u = 68 please comment if you have any doubts will clarify